OpenStudy (anonymous):

x-9=-6x+5 WILL GIVE MEDAL FOR CORRECT ANSWER!

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

5+9=14 and 6x+x=7x 7x=14 x=2

OpenStudy (anonymous):

Add 6x to both sides: 7x-9=5 Add nine to both sides: 7x=14 Divide by seven: x=2\[x=2\]

OpenStudy (anonymous):

thanks :P

OpenStudy (anonymous):

but yours was the best

OpenStudy (anonymous):

:) It doesn't cost us anything to give each other medals.

OpenStudy (anonymous):

well i like medals! :)

OpenStudy (anonymous):

haha